OpenStudy (anonymous):

sketch a graph of y=xpower of 2 if the graph is translated 1 unit to the left, write down the equation of the new graph (3)marks

OpenStudy (anonymous):

You know there is a vertex at x=0. If you shift the graph 1 unit to the left, what does x have to be to produce a vertex at -1?

OpenStudy (anonymous):

NEW EQN OF GRAPH IS (1-x^2)=y

OpenStudy (anonymous):

Nope...

OpenStudy (anonymous):

you must first draw a table from negetive 3 to positive 3 and in the place of x you must put those numbers e.g. y=x^2 therefore -3^2 -3 is in (x)'s place.

OpenStudy (anonymous):

you still there bra

OpenStudy (anonymous):

See y = x^2  Now plot graph for this and you will get  Now what will happen when you move this graph to the left by 1 unit ??